Help Please !
The hortizontal sides of a rectangle are one half of the length of the sides of a separate square. the vertical sides of the rectangle are 4 cm more than three times the side of a square. if the perimeter of the rectangle is twice the perimeter of the square what are the dimensions of the rectangle?

Answer by chessace(471)

Asked to find h and v, let s=side of square.
Translate english to algebra:
h=0.5*s
v=4cm+3*s
2(h+v)=2*(4s)
Now probably easiest to solve for s first:
8s=2(0.5s+4+3s)
8s=7s+8
s=8cm
h=4cm
v=28cm